OpenStudy (darkbeautystar):

Juan deposited $200 in a savings account earning 3% interest over 2 years. What was the total amount of interest earned at the end of 2 years? A. $12 B. $24 C. $120 D. $240

OpenStudy (darkbeautystar):

@johnweldon1993

OpenStudy (darkbeautystar):

I just need help checking my answers..I have a five question test and I have gotten all of the answers ready

OpenStudy (johnweldon1993):

Simple interest \[\large I = PRT\] The interest accrued would be \(\large I = 200*0.03*2 = ?\)

OpenStudy (darkbeautystar):

12-Thats what I had

OpenStudy (darkbeautystar):

@johnweldon1993 - I got $3600 Mai borrowed $12,000 at 5% interest for 6 years. What was the total interest Mai paid?

OpenStudy (johnweldon1993):

\[\large \color \red{\checkmark}\]

OpenStudy (darkbeautystar):

YAY

OpenStudy (darkbeautystar):

@johnweldon1993 For this question I only found T..and I don't even think that it is right! 2.833333332 - That is T. (Supposedly) Katie invested $1200 at 8.5% interest. She earned $612 interest at the end of y years. What is the value of y?

OpenStudy (darkbeautystar):

Sorry I thought I had this one

OpenStudy (darkbeautystar):

Actually is it 6

OpenStudy (johnweldon1993):

\[\large I = PRT\] \[\large T = \frac{I}{PR} = \frac{612}{1200(0.085)} = 6\] There ya go :)

OpenStudy (darkbeautystar):

@johnweldon1993 - I think this one is 6 again! Lily invested $750 in a savings account and earned $225 in interest at the end of 5 years. What was the interest rate? A. 5% B. 5.5% C. 6% D. 6.5%

OpenStudy (darkbeautystar):

Thanks so much for your help!

OpenStudy (johnweldon1993):

You're right again! Great job! :)

OpenStudy (darkbeautystar):

Thanks

OpenStudy (darkbeautystar):

Do you think you could help me with ratios

OpenStudy (darkbeautystar):

I am terrible at ratios

OpenStudy (darkbeautystar):

I will open a new question